OpenStudy (anonymous):

What belongs in the box marked with the question mark in the proof?Given: 3x - 4 = 2(x + 8), Prove: x = 20

OpenStudy (anonymous):

http://prntscr.com/cojcl here is the image.

OpenStudy (anonymous):

@SmoothMath can you help?

OpenStudy (anonymous):

Would it be The Distributive Property?

OpenStudy (anonymous):

adding 4 on both sides

OpenStudy (anonymous):

This is a Two-Column proof, the reasons have to be postulate, theorems, properties, etc.

OpenStudy (anonymous):

@zepp I've heard you're good with two column proofs

OpenStudy (zepp):

We've send 2x to the left side, I believe it should be Associative Property of Addition 3x = 2x + 20 3x + 0 = 2x + 20

OpenStudy (anonymous):

Zepp, the question mark is in the column after the statement 3x = 2x + 20 so it needs to be a reason for going from 3x-4 = 2x + 20 to 3x = 2x+20 You should ask yourself how it is that we're able to get from one to the other, and the answer, as rabhajishek point out, is that we added 4 to both sides. So that's what we DID, but what is it that ALLOWS us to do that? The associative property of addition isn't it, I wouldn't say. That's the one that would allow us to rewrite (a+b)+c as a+(b+c) Instead, there's a property that lets us add something to both sides of an equation without unbalancing it. What's that property called? The addition property of equality.

OpenStudy (anonymous):

Sorry for the long response time, Feara. I was nappin' :3

OpenStudy (zepp):

Oh I see :D

OpenStudy (zepp):

Thanks for the clarification @SmoothMath